Operation Summary for April 7, 2010 - April 9, 2010
Wednesday April 7
The day shift began with Operations monitoring the Tevatron (TeV) store 7730 (86.49E30), the Recycler stash (251.0E10), and the Antiproton Source stack (17.0E10), with the Linac, Booster, and Main Injector operating normally, with MiniBooNE off due to a IRM crate trigger problem, with the NuMI and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, MiniBooNE started taking beam by 10:49 AM after their controls issues were resolved. Operations terminated store 7730 at 2:37 PM and put the TeV into HEP shot setup. Operations established store 7731 at 3:52 PM with an average initial luminosity (AIL) of 323.5E30. An MTest corrector magnet power supply blew a fuse causing a short downtime. Operators continued Booster extraction tuning all shift.
During the evening shift, Operations disabled beam to MTest at 6:04 PM. A controls CAMAC crate power supply failed, tripping off a Recycler bulk power supply, which caused the loss of 145E10 antiprotons from the Recycler. Operators replaced the power supply and beam operations resumed.
Thursday April 8
The midnight shift began with Operations monitoring store 7731 (109.3E30) and the stash (51.0E10), with Pbar stacking (11.8E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to MTest at 4:28 AM. A large helium compressor at CHL tripped off, which caused a loss of pressure in the helium transfer lines. CHL personnel were able to start a backup compressor. The transfer line pressures were near normal by the end of shift.During the day shift, Tevatron experts conducted studies for the last two hours of store 7731. Operations terminated store 7731 at 2:09 PM and put the TeV into HEP shot setup. Operations established store 7732 at 3:21 PM with an initial luminosity of 324.4E30. A Main Injector low level RF reboot caused some controls issues that took about 30 minutes to resolve. During the evening shift, Store 7732 collided all shift. Operations disabled beam to MTest at 6:00 PM. The I- source vacuum valve closed causing a few minutes of downtime. Extensive Main Injector tuning was performed to reduce losses.
Friday April 9
The midnight shift began with Operations monitoring store 7732 (103.3E30) and the stash (245.1E10), with Pbar stacking (6.3E10), and with NuMI and MiniBooNE taking beam.During the midnight shift, Operations performed Booster and Main Injector tuning to reduce losses. Operations and a system expert accessed the H- source to repair a controls issue, this caused minimal downtime. Operations established beam to M-Test at 4:27 AM. Operations terminated store 7732 at 7:52 AM.
The Plans for Friday and the Weekend
The plan for today is to put in a TeV store, then stack, stash, and store through the weekend.
Comments and Suggestions
|last modified 04/09/2010 email Fermilab|
|Security, Privacy, Legal|